OpenStudy (anonymous):

find the derivative during delta/first principl method of f(r) = pie(r)^2

OpenStudy (anonymous):

i'm not sure exactly what the delta/first principle method is, but the derivative of pie r squared is \[2\pi r\]

OpenStudy (anonymous):

that i know its 2pie r I needed to do it in this way y= pr^2 p= pie y+ dy= p(r+dr)^2 now we need do find dy/dr, hot to proceed

OpenStudy (anonymous):

\[\frac{dy}{dr}=\lim_{h \rightarrow 0} \pi \frac{((r+h)^2 - r^2)}{h}\]\[\lim_{h \rightarrow 0} \pi \frac{((r^2+2rh+h^2) - r^2)}{h}\]\[\lim_{h \rightarrow 0} \pi \frac{h(2r+h)}{h}\]\[\lim_{h \rightarrow 0} \pi (2r+h)\]\[ \pi (2r+0)\]\[ \pi (2r)\] Sorry, I've missed all the equal signs

OpenStudy (anonymous):

Sorry, but can you explain?

OpenStudy (anonymous):

I hope this would help. I thought that was the latex problem.Sorry again if I've brought you some troubles

OpenStudy (anonymous):

i'm not quite sure what you're asking, i'm sorry :( :(

OpenStudy (anonymous):

I GOT IT THANKS